OpenStudy (anonymous):

Find f o g if f(x) = 2x - 1 and g(x) = x + 2.

OpenStudy (anonymous):

f o g (x) = 2(x + 2)-1 = 2x+4-1 = 2x+3

OpenStudy (anonymous):

If y varies inversely as x. find the constant of variation if y = 6 and x = 2?